# Beyond Fast and Slow: Cognitive-Inspired El-astic Reasoning for Large Language Models

### v1

@@ A.2 Large Reasoning Models
Large language models (LLMs) have demonstrated remarkable capabilities in natural language understanding and complex reasoning (grattafiori2024llama), becoming a pivotal advancement in AI. To further enhance performance in ”System 2” reasoning domains (li2025system) such as mathematics (cobbe2021training; hendrycks2021measuring) and programming (chen2021evaluating), which require deep thought, researchers have developed specialized Large Reasoning Models (LRMs) (xu2025towards).
Fundamentals of Large Reasoning Models. The core of LRMs lies in internalizing and enhancing Chain-of-Thought (CoT) (wei2022chain) reasoning through strategies like Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). CoT significantly improves LLM performance on complex tasks by prompting the model to generate a series of intermediate reasoning steps, with variants like Self-Consistency CoT (wang2022self), Tree-of-Thoughts (yao2023tree), and Graph-of-Thoughts (besta2024graph) emerging. LRMs aim to integrate this step-by-step reasoning capability more deeply within the model, rather than relying solely on explicit test-time prompts or external augmentations, by generating detailed, structured reasoning sequences to achieve higher accuracy. Prominent examples of LRMs include OpenAI’s o1 model and DeepSeekAI’s DeepSeek-R1 model.
Training Mechanisms of LRMs. The training mechanisms for LRMs typically combine SFT to learn from high-quality reasoning paths and RL to further optimize reasoning strategies, enabling exploration of better problem-solving steps (luo2025o1; aggarwal2025l1). For instance, DeepSeek-R1 enhanced its general reasoning capabilities through multiple rounds of SFT and RL, emphasizing structured thinking templates and rule-based reward mechanisms. OpenAI’s o1 is speculated by the community to employ tree-search methods like Monte Carlo Tree Search (MCTS) (coulom2006efficient) combined with a Process Reward Model (PRM) (uesato2022solving) to explore and evaluate different reasoning paths. These advanced training methods enable LRMs to generate complex thought processes internally, progressively deriving final answers, demonstrating significant potential in solving challenging mathematical problems and programming tasks, as assessed by benchmarks like Sys2Bench (parashar2025inference).
@@ A.3 Reinforcement Learning
Reinforcement Learning (RL) (kaelbling1996reinforcement; hucontinual; sun2025open) constitutes a paradigm within machine learning wherein an agent learns to optimize its decision-making process through interaction with an environment. This interaction involves performing actions and receiving consequent feedback, typically in the form of rewards or penalties. The principal learning objective in RL is the maximization of a cumulative reward signal. In contrast to supervised learning, which relies on datasets comprising pre-defined input-output pairs for model training, RL entails an agent acquiring knowledge from the repercussions of its actions, mediated by this reward-penalty mechanism. This iterative, trial-and-error learning process, coupled with its emphasis on sequential decision-making under uncertainty, distinguishes RL from supervised learning methodologies that depend on labeled datasets. Existing reinforcement learning algorithms can be broadly categorized based on whether an explicit model of the environment is learned or utilized, leading to two principal classes: Model-free RL and Model-based RL.
@@
Model-free RL. Model-free RL algorithms enable the agent to learn optimal policies directly from trajectory samples accrued through interaction with the environment, without explicitly constructing an environmental model. Within model-free RL, algorithms are further distinguished by the components they learn, leading to three primary sub-categories: actor-only, critic-only, and actor-critic algorithms. Actor-only algorithms directly learn a policy network, denoted as $\pi_{\theta}(a|s)$, which maps states to actions. This network takes the current state $s_{t}$ as input and outputs the action $a_{t}$. Prominent examples of such algorithms include Reinforce (williams1992simple) and various policy gradient methods (sutton1999policy). Critic-only algorithms, in contrast, focus solely on learning a value function (e.g., state-value or action-value function). Given a state $s_{t}$, the learned value model is used to evaluate all possible actions $a^{\prime}\in A$, and the action $a_{t}$ yielding the maximum estimated value is selected. This category encompasses methods such as Q-learning (watkins1989learning). Actor-critic algorithms combine these two approaches by concurrently maintaining and learning both a policy network (the actor) for action selection and a value function model (the critic) for evaluating actions or states. This category includes algorithms such as Deep Deterministic Policy Gradient (DDPG) (lillicrap2015continuous), Trust Region Policy Optimization (TRPO) (schulman2015trust), Proximal Policy Optimization (PPO) (ppo), and Asynchronous Advantage Actor-Critic (A3C) (mnih2016asynchronous). Notably, PPO has gained considerable traction for training large language models. Recent advancements in this area include GRPO (grpo2024), which employs group‑based advantage estimates within a KL‑regularized loss function to reduce computational overhead and enhance update stability, and DAPO (dapo2024), which utilizes distinct clipping mechanisms and adaptive sampling techniques to improve efficiency and reproducibility during the fine‑tuning of large‑scale models.
@@
Model-based RL. Model-based RL algorithms endeavor to learn an explicit model of the environment, thereby addressing challenges related to sample efficiency. This is because the agent can leverage the learned model for planning and decision-making, reducing the necessity for extensive direct environmental interaction. The learned representation of the environment is commonly termed a ’world model’. This world model typically predicts the subsequent state $s_{t+1}$ and the immediate reward $r_{t}$ based on the current state $s_{t}$ and the action $a_{t}$ taken. Exemplary model-based RL algorithms include Dyna-Q (peng2018deep), Model-Based Policy Optimization (MBPO) (janner2019trust), and Adaptation Augmented Model-based Policy Optimization (AMPO) (shen2023adaptation).
@@ A.4 Tool Integrated Reasoning
Foundations and Evaluation of Tool-Integrated Reasoning. Early research predominantly focused on equipping LLMs with external tools to overcome their inherent limitations. This involved introducing concepts such as program executors (chen2022program) and search engines (vu2023freshllms) to enhance their problem-solving capabilities (qin2024tool). The core tenet of TIR is to enable LLMs to interact with these external tools, thereby addressing issues such as outdated knowledge, computational inaccuracies, and shallow reasoning (qian2025toolrl). As research progressed, a series of specialized benchmarks were proposed to systematically evaluate model performance in tool selection, argument generation, and generalization (qin2023toolllm). Concurrently, the construction of high-quality tool-use datasets became a significant driver for advancements in the field (liu2024toolace; qian2025smart), and these datasets and benchmarks have further facilitated the exploration of TIR techniques across diverse modalities and specialized domains (shen2025vlm).
@@
Supervised Fine-Tuning for Tool-Integrated Reasoning. In the initial stages of training LLMs for TIR tasks, Supervised Fine-Tuning (SFT) was the predominant approach. These methods typically relied on offline-generated tool-use trajectories, upon which models were subsequently fine-tuned (chen2024agent; acikgoz2025can). For instance, in code-integrated reasoning scenarios, researchers endowed models with initial capabilities for code invocation and result interpretation by performing SFT on self-curated Chain-of-Thought (CoT) data that included code execution steps (chen2025empirical). However, SFT methods exhibit notable deficiencies in terms of generalization, exploration, and adaptability (chu2025sft; guo2025deepseek). Models often merely imitate specific patterns within the training data, struggling to adapt to unseen or more complex tool-use scenarios and failing to autonomously learn when and how to invoke external tools most effectively (feng2025retool).
@@
Reinforcement Learning for Tool-Integrated Reasoning. To overcome the limitations inherent in SFT, Reinforcement Learning (RL) has emerged as a promising paradigm for training more adaptive and generalizable tool-using LLMs. RL frameworks enable models to learn optimal tool invocation strategies through direct interaction with environments and feedback signals, moving beyond the mere imitation of static trajectories. Various RL algorithms, including Proximal Policy Optimization (PPO) and Direct Preference Optimization (DPO) , are being adapted to the specific challenges of TIR. Central to applying RL in TIR is the development of effective feedback mechanisms. One stream of research focuses on meticulous reward engineering, designing rewards that offer step-grained guidance on tool invocation correctness and contribution (yu2024steptool), or penalize specific error types (ye2024tl). Another prominent trend involves learning from preferences or comparative feedback, often leveraging techniques like Direct Preference Optimization (DPO) or ranking losses. This allows models to learn from a broader spectrum of execution traces, including imperfect or erroneous paths, by comparing preferred outcomes against less desirable ones (chen2024advancing; zeng2025boosting; jung2025diatool). Such approaches also extend to optimizing multi-turn dialogue control and learning from varied forms of execution feedback (wu2024toolplanner). Collectively, these RL strategies aim to enhance not only the precision of tool selection and usage but also the model’s nuanced decision-making in complex, interactive scenarios.
@@ Appendix C Benchmarks
  • GSM8K (cobbe2021training) is a widely adopted benchmark for evaluating the arithmetic reasoning capabilities of language models on grade-school math word problems. The dataset comprises 7,473 training examples and 1,319 test examples. Problems are designed to necessitate 2 to 8 steps of chain-of-thought reasoning for their solution. These problems are human-curated to ensure linguistic diversity and assess the model’s understanding of fundamental mathematical concepts and its ability to perform multi-step arithmetic operations. The dataset is available at this link and under the MIT License.
  • MATH (hendrycks2021measuring), is a challenging competition-level mathematics dataset designed to measure reasoning capabilities in advanced mathematics. It encompasses problems from pre-algebra, algebra, geometry, number theory, and calculus, among others, totaling 7,500 training examples and 5,000 test examples. These problems typically require complex symbolic manipulation, abstract thinking, and multi-step deductive reasoning. The dataset is available at this link and under the MIT License.
@@
  • MAWPS (koncel2016mawps) is a benchmark focusing on fundamental math word problems (MWPs). It aggregates problems from various sources, primarily involving one or a few arithmetic steps, with a difficulty level roughly corresponding to elementary school mathematics. MAWPS contains 238 test instances and is designed to evaluate model robustness to variations in problem phrasing and numerical values. This dataset is under the MIT License and can be found within the MWPToolkit project at this link.
  • CollegeMath (tang2024mathscale) is an emerging dataset designed to assess reasoning abilities on university-level mathematics problems. It comprises 2,818 problems meticulously curated from 9 different university mathematics textbooks, spanning seven core areas such as linear algebra, calculus, probability theory, and differential equations. CollegeMath problems test not only computational skills but also the understanding of advanced mathematical concepts, abstract reasoning, and the application of theorems and methods in complex scenarios, thereby posing a significant challenge to models’ generalization and deep reasoning capabilities. This publicly available dataset serves as a valuable resource for evaluating LLMs in the domain of higher education mathematics and is available at this link. For our evaluation, we used a randomly sampled subset of 1,200 items from this dataset.
  • CommonsenseQA (talmor2019commonsenseqa) is a multiple-choice question-answering dataset designed to evaluate commonsense reasoning. It contains 12,247 questions, each with 5 options. The model is required to select the most plausible answer, a task that typically cannot be resolved by simple keyword matching but necessitates an understanding of world knowledge, conceptual relationships, and implicit information. CommonsenseQA challenges models’ reasoning abilities in non-formal and everyday contexts. The dataset is available at this link and under the CC-BY-4.0 License.
  • MedQA (jin2021disease) is a professional medical question-answering dataset, with content derived from licensing examinations in the United States (USMLE), mainland China (NMLEC), and Taiwan (TMQE). Presented primarily as multiple-choice questions, it covers a broad spectrum of medical subfields, including clinical medicine, basic sciences, pharmacology, and diagnostics. MedQA aims to assess models’ knowledge mastery, information retrieval capabilities, and, to some extent, clinical reasoning within a highly specialized domain. This dataset is crucial for advancing research into LLM applications in critical sectors like healthcare. For our experiments, we specifically utilized the official ”US” test split from the MedQA dataset. This portion consists of multiple-choice questions designed to evaluate the final test performance on US medical examination questions. The dataset is available at this link and under the MIT License.

### v2

@@ A.2 Large Reasoning Models
Large language models (LLMs) have demonstrated remarkable capabilities in natural language understanding and complex reasoning (Team, 2024), becoming a pivotal advancement in AI. To further enhance performance in ”System 2” reasoning domains (Li et al., 2025) such as mathematics (Cobbe et al., 2021; Hendrycks et al., 2021) and programming (Chen et al., 2021), which require deep thought, researchers have developed specialized Large Reasoning Models (LRMs) (Xu et al., 2025a).
Fundamentals of Large Reasoning Models. The core of LRMs lies in internalizing and enhancing Chain-of-Thought (CoT) (Wei et al., 2022b) reasoning through strategies like Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL). CoT significantly improves LLM performance on complex tasks by prompting the model to generate a series of intermediate reasoning steps, with variants like Self-Consistency CoT (Wang et al., 2023), Tree-of-Thoughts (Yao et al., 2023), and Graph-of-Thoughts (Besta et al., 2024) emerging. LRMs aim to integrate this step-by-step reasoning capability more deeply within the model, rather than relying solely on explicit test-time prompts or external augmentations, by generating detailed, structured reasoning sequences to achieve higher accuracy. Prominent examples of LRMs include OpenAI’s o1 model and DeepSeekAI’s DeepSeek-R1 model.
Training Mechanisms of LRMs. The training mechanisms for LRMs typically combine SFT to learn from high-quality reasoning paths and RL to further optimize reasoning strategies, enabling exploration of better problem-solving steps (Luo et al., 2025; Aggarwal and Welleck, 2025). For instance, DeepSeek-R1 enhanced its general reasoning capabilities through multiple rounds of SFT and RL, emphasizing structured thinking templates and rule-based reward mechanisms. OpenAI’s o1 is speculated by the community to employ tree-search methods like Monte Carlo Tree Search (MCTS) (Coulom, 2006) combined with a Process Reward Model (PRM) (Uesato et al., 2022) to explore and evaluate different reasoning paths. These advanced training methods enable LRMs to generate complex thought processes internally, progressively deriving final answers, demonstrating significant potential in solving challenging mathematical problems and programming tasks, as assessed by benchmarks like Sys2Bench (Parashar et al., 2025).
@@ A.3 Reinforcement Learning
Reinforcement Learning (RL) (Kaelbling et al., 1996) constitutes a paradigm within machine learning wherein an agent learns to optimize its decision-making process through interaction with an environment. This interaction involves performing actions and receiving consequent feedback, typically in the form of rewards or penalties. The principal learning objective in RL is the maximization of a cumulative reward signal. In contrast to supervised learning, which relies on datasets comprising pre-defined input-output pairs for model training, RL entails an agent acquiring knowledge from the repercussions of its actions, mediated by this reward-penalty mechanism. This iterative, trial-and-error learning process, coupled with its emphasis on sequential decision-making under uncertainty, distinguishes RL from supervised learning methodologies that depend on labeled datasets. Existing reinforcement learning algorithms can be broadly categorized based on whether an explicit model of the environment is learned or utilized, leading to two principal classes: Model-free RL and Model-based RL.
@@
Model-free RL. Model-free RL algorithms enable the agent to learn optimal policies directly from trajectory samples accrued through interaction with the environment, without explicitly constructing an environmental model. Within model-free RL, algorithms are further distinguished by the components they learn, leading to three primary sub-categories: actor-only, critic-only, and actor-critic algorithms. Actor-only algorithms directly learn a policy network, denoted as $\pi_{\theta}(a|s)$, which maps states to actions. This network takes the current state $s_{t}$ as input and outputs the action $a_{t}$. Prominent examples of such algorithms include Reinforce (Williams, 1992) and various policy gradient methods (Sutton et al., 1999). Critic-only algorithms, in contrast, focus solely on learning a value function (e.g., state-value or action-value function). Given a state $s_{t}$, the learned value model is used to evaluate all possible actions $a^{\prime}\in A$, and the action $a_{t}$ yielding the maximum estimated value is selected. This category encompasses methods such as Q-learning (WATKINS, 1989). Actor-critic algorithms combine these two approaches by concurrently maintaining and learning both a policy network (the actor) for action selection and a value function model (the critic) for evaluating actions or states. This category includes algorithms such as Deep Deterministic Policy Gradient (DDPG) (Tkachenko, 2015), Trust Region Policy Optimization (TRPO) (Schulman et al., 2015), Proximal Policy Optimization (PPO) (Schulman et al., 2017), and Asynchronous Advantage Actor-Critic (A3C) (Mnih et al., 2016). Notably, PPO has gained considerable traction for training large language models. Recent advancements in this area include GRPO (Shao et al., 2024), which employs group‑based advantage estimates within a loss function to reduce computational overhead and enhance update stability, and DAPO (Yu et al., 2025), which utilizes distinct clipping mechanisms and adaptive sampling techniques to improve efficiency and reproducibility during the fine‑tuning of large‑scale models.
@@
Model-based RL. Model-based RL algorithms endeavor to learn an explicit model of the environment, thereby addressing challenges related to sample efficiency. This is because the agent can leverage the learned model for planning and decision-making, reducing the necessity for extensive direct environmental interaction. The learned representation of the environment is commonly termed a ’world model’. This world model typically predicts the subsequent state $s_{t+1}$ and the immediate reward $r_{t}$ based on the current state $s_{t}$ and the action $a_{t}$ taken. Exemplary model-based RL algorithms include Dyna-Q (Peng et al., 2018), Model-Based Policy Optimization (MBPO) (Janner et al., 2019), and Adaptation Augmented Model-based Policy Optimization (AMPO) (Shen et al., 2023).
@@ A.4 Tool Integrated Reasoning
Foundations and Evaluation of Tool-Integrated Reasoning. Early research predominantly focused on equipping LLMs with external tools to overcome their inherent limitations. This involved introducing concepts such as program executors (Chen et al., 2023) and search engines (Vu et al., 2024) to enhance their problem-solving capabilities (Qin et al., 2025). The core tenet of TIR is to enable LLMs to interact with these external tools, thereby addressing issues such as outdated knowledge, computational inaccuracies, and shallow reasoning (Qian et al., 2025a). As research progressed, a series of specialized benchmarks were proposed to systematically evaluate model performance in tool selection, argument generation, and generalization (Qin et al., 2024). Concurrently, the construction of high-quality tool-use datasets became a significant driver for advancements in the field (Liu et al., 2025; Qian et al., 2025b), and these datasets and benchmarks have further facilitated the exploration of TIR techniques across diverse modalities and specialized domains (Shen et al., 2025).
@@
Supervised Fine-Tuning for Tool-Integrated Reasoning. In the initial stages of training LLMs for TIR tasks, Supervised Fine-Tuning (SFT) was the predominant approach. These methods typically relied on offline-generated tool-use trajectories, upon which models were subsequently fine-tuned (Chen et al., 2024b; Acikgoz et al., 2025). For instance, in code-integrated reasoning scenarios, researchers endowed models with initial capabilities for code invocation and result interpretation by performing SFT on self-curated Chain-of-Thought (CoT) data that included code execution steps (Chen et al., 2025b). However, SFT methods exhibit notable deficiencies in terms of generalization, exploration, and adaptability (Chu et al., 2025; DeepSeek-AI, 2025). Models often merely imitate specific patterns within the training data, struggling to adapt to unseen or more complex tool-use scenarios and failing to autonomously learn when and how to invoke external tools most effectively (Feng et al., 2025).
@@
Reinforcement Learning for Tool-Integrated Reasoning. To overcome the limitations inherent in SFT, Reinforcement Learning (RL) has emerged as a promising paradigm for training more adaptive and generalizable tool-using LLMs. RL frameworks enable models to learn optimal tool invocation strategies through direct interaction with environments and feedback signals, moving beyond the mere imitation of static trajectories. Various RL algorithms, including Proximal Policy Optimization (PPO) and Direct Preference Optimization (DPO) , are being adapted to the specific challenges of TIR. Central to applying RL in TIR is the development of effective feedback mechanisms. One stream of research focuses on meticulous reward engineering, designing rewards that offer step-grained guidance on tool invocation correctness and contribution (Yu et al., 2024), or penalize specific error types (Ye et al., 2024). Another prominent trend involves learning from preferences or comparative feedback, often leveraging techniques like Direct Preference Optimization (DPO) or ranking losses. This allows models to learn from a broader spectrum of execution traces, including imperfect or erroneous paths, by comparing preferred outcomes against less desirable ones (Chen et al., 2024a; Zeng et al., 2025; Jung et al., 2025). Such approaches also extend to optimizing multi-turn dialogue control and learning from varied forms of execution feedback (Wu et al., 2024). Collectively, these RL strategies aim to enhance not only the precision of tool selection and usage but also the model’s nuanced decision-making in complex, interactive scenarios.
@@ Appendix C Benchmarks
  • GSM8K (Cobbe et al., 2021) is a widely adopted benchmark for evaluating the arithmetic reasoning capabilities of language models on grade-school math word problems. The dataset comprises 7,473 training examples and 1,319 test examples. Problems are designed to necessitate 2 to 8 steps of chain-of-thought reasoning for their solution. These problems are human-curated to ensure linguistic diversity and assess the model’s understanding of fundamental mathematical concepts and its ability to perform multi-step arithmetic operations. The dataset is available at this link and under the MIT License.
  • MATH (Hendrycks et al., 2021), is a challenging competition-level mathematics dataset designed to measure reasoning capabilities in advanced mathematics. It encompasses problems from pre-algebra, algebra, geometry, number theory, and calculus, among others, totaling 7,500 training examples and 5,000 test examples. These problems typically require complex symbolic manipulation, abstract thinking, and multi-step deductive reasoning. The dataset is available at this link and under the MIT License.
@@
  • MAWPS (Koncel-Kedziorski et al., 2016) is a benchmark focusing on fundamental math word problems (MWPs). It aggregates problems from various sources, primarily involving one or a few arithmetic steps, with a difficulty level roughly corresponding to elementary school mathematics. MAWPS contains 238 test instances and is designed to evaluate model robustness to variations in problem phrasing and numerical values. This dataset is under the MIT License and can be found within the MWPToolkit project at this link.
  • CollegeMath (Tang et al., 2024) is an emerging dataset designed to assess reasoning abilities on university-level mathematics problems. It comprises 2,818 problems meticulously curated from 9 different university mathematics textbooks, spanning seven core areas such as linear algebra, calculus, probability theory, and differential equations. CollegeMath problems test not only computational skills but also the understanding of advanced mathematical concepts, abstract reasoning, and the application of theorems and methods in complex scenarios, thereby posing a significant challenge to models’ generalization and deep reasoning capabilities. This publicly available dataset serves as a valuable resource for evaluating LLMs in the domain of higher education mathematics and is available at this link. For our evaluation, we used a randomly sampled subset of 1,200 items from this dataset.
  • CommonsenseQA (Talmor et al., 2019) is a multiple-choice question-answering dataset designed to evaluate commonsense reasoning. It contains 12,247 questions, each with 5 options. The model is required to select the most plausible answer, a task that typically cannot be resolved by simple keyword matching but necessitates an understanding of world knowledge, conceptual relationships, and implicit information. CommonsenseQA challenges models’ reasoning abilities in non-formal and everyday contexts. The dataset is available at this link and under the CC-BY-4.0 License.
  • MedQA (Jin et al., 2020) is a professional medical question-answering dataset, with content derived from licensing examinations in the United States (USMLE), mainland China (NMLEC), and Taiwan (TMQE). Presented primarily as multiple-choice questions, it covers a broad spectrum of medical subfields, including clinical medicine, basic sciences, pharmacology, and diagnostics. MedQA aims to assess models’ knowledge mastery, information retrieval capabilities, and, to some extent, clinical reasoning within a highly specialized domain. This dataset is crucial for advancing research into LLM applications in critical sectors like healthcare. For our experiments, we specifically utilized the official ”US” test split from the MedQA dataset. This portion consists of multiple-choice questions designed to evaluate the final test performance on US medical examination questions. The dataset is available at this link and under the MIT License.